OpenStudy (anonymous):

which statement best explains how the value of y changes each time x is increased by 1 unit? (x/3)-(y/2)=-1

OpenStudy (anonymous):

can anybody help me with this one

OpenStudy (anonymous):

HELP!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

"which statement best explains" Where are the statements ?

OpenStudy (anonymous):

first solve for y: \[\frac{x}{3} - \frac{y}{2} = -1 \;\;\;\;|-\frac{x}{3}\]\[-\frac{y}{2} = -\frac{x}{3} - 1 \;\;\;\; | *(-2) \]\[y = \frac{2}{3}x + 2\] Then lets say we look at an arbitrary x. Then y is equal to the equation above. Then look at what y becomes if we substitute x+1 for x in this equation. \[y _{x} = \frac{2}{3}x + 2\]\[y _{x+1} = \frac{2}{3}(x+1) + 2 = \frac{2}{3}x+\frac{2}{3} + 2 = \frac{2}{3}x + 2 + \frac{2}{3}\] If you don't see it yet you can do a subtraction and find that \[y_{x+1} - y_{x} = \frac{2}{3} \implies y_{x+1} = y_{x} + \frac{2}{3}\]And that is the statement we were looking for, explicitly describing the new y value we get when increasing x by 1.